You serve the tennis ball to your opponent from a hight is 2.6m with an initial velocity of 20.4m/s. What is the time it takes to reach your opponent? What is the total distance it travels? What is the final vertical velocity? What is the final velocity?
1
Expert's answer
2019-12-19T11:36:40-0500

Time:


"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2(2.6)}{9.8}}=0.728\ s"

The total distance:


"d=v_xt=(20.4)(0.728)=15\ m"

The final vertical velocity:


"v_y=\sqrt{2gh}=\sqrt{2(9.8)(2.6)}=7.1\frac{m}{s}"

The final velocity:


"v=\sqrt{7.1^2+20.4^2}=21.6\frac{m}{s}"